VEGAS Team Meeting

Dial in Number: 877-937-8961

Passcode: 9445091#

Attendees: Mark, Richard, Ray, Randy, Adam, John, Amanda, Sam, Mark, Paul, Anish, Marty, Jason

Agenda:

Whole group:

  1. Weird VEGAS bug (see attached plot)

GB only

  1. Release of last BOF file
  2. Status of Mode 10-19 firmware fix
  3. Status of commissioning observations
  4. Review of problems from: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCommissioningLogs
  5. Status of Jay Lockman Tsys bug
  6. Review of DIBAS upgrade memo
  7. New version of l8 (l8ver148)
  8. Test Plans

Meeting notes:

Whole group:

  1. Weird VEGAS bug (see attached plot)
    • low priority if this is a one-off result, but if this is a regular result this is a big problem.
    • Continue to monitor for reoccurance

GB only

  1. Release of last BOF file
    • all have been built and calibrated
    • Test and release
  2. Status of Mode 10-19 firmware fix
    • Released new version of l8 (l 8ver148)
    • SRBS tests seem to work, but the quality of spectra on SRBS is not really that good
    • Try a POC test using VEGAS's internal tone generator, an external tone generator, or our Cross-Coupling Test Fixture as our signal source - needs a SDD person to commission
      • The software fixes must be removed before tests can proceed
  3. Status of commissioning observations
    • a large number of scans are in the data analysis mode
    • So far no anomalies
    • Ripple tests will need to be redone - Amanda will analyze data
  4. Review of problems from: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCommissioningLogs
  5. Status of Jay Lockman Tsys bug
    • one integration out of many have a different Tsys
    • Maybe first/last scan?
    • Will work with Jay's data or re-run to try and identify problem
  6. Review of DIBAS upgrade memo
    • Need to quantify "slight change of functionality" comment
    • Please confirm it can be sent
    • Means GB needs to be ready to support SHAO after they receive upgrades
      • SHAO to propose when - GB can delay if other VEGAS work is occupying staff
  7. New version of l8 (l 8ver148)
  8. Test plans for this week
    • Unscheduled request(s)
      • Test remaining BOF files (Richard & Adam to schedule)
      • Ripple tests (sky, but not movement) (Richard & Adam to schedule)
      • Test firmware fix for Modes 10-19 (Randy & Randy to schedule)
    • Tuesday
    • Wednesday
    • Thursday
    • Monday (8/4)
      • None at this time

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-07-28
Topic revision: r2 - 2015-01-27, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback